# Concatenate, Fine-tuning, Re-training: A SAM-enabled Framework for Semi-supervised 3D Medical Image Segmentation

### v1

@@ VIII-D1 Single Target Segmentation
We evaluate the impact of CFR with scarce annotations on the LA and BraTS datasets. In Table III, when only one labeled volume is available, supervised learning performs poorly, achieving only 17.99% accuracy. CFR takes advantage of medical image concatenation and SAM’s powerful feature capture capabilities to guide MT to increase from 29.68% to 74.40%. The results of the BraTS are shown in Table IV and CFR helps MT improve by 15.06% and MagicNet by 15.15%. In addition, one phenomenon observed is that some semi-supervised methods perform worse than the lower bound. This could be attributed to the incomplete similarity of the data distribution between labeled and unlabeled samples, and the inaccurate pseudo-labels from the unlabeled samples have a negative impact on the model’s learning process.
VIII-D2 Multiple Target Segmentation
@@
The results of scarce annotations on two multi-organ datasets are presented in Table VII and Table VIII. On BTCV dataset with only one labeled sample, the model struggles to identify esophagus, aorta, inferior vena cava, portal and splenic veins, left and right adrenal glands, with Dice scores below 5% for left kidney, stomach, and pancreas. Semi-supervised methods with unlabeled samples alleviate the poor performance of most organs. Our CFR framework achieves improvements of 10.12% and 13.57% on MT and MagicNet models, respectively, particularly with a nearly 30% improvement on aorta region. Similarly, on MACT dataset, CFR helps identify challenging classes, with significant improvements of 37.33% for spleen and 34.53% for right kidney on the Mean Teacher model.
VIII-E Effectiveness of Re-training Module
We report the results of fine-tuning step omitting re-training in all datasets in Table IX, and compare the model parameters and performance after fine-tuning and re-training steps. By comparing fine-tuning results with re-training results in more detail, we observe that there is a slight performance decrease in CFR with scarce labels on LA and BTCV datasets, but overall there is a clear improvement in effect. On the other hand, considering the practical deployment requirements, re-training model has a smaller number of parameters than fine-tuning model. Our fine-tuning module CFR has the same parameter size as the foundation model (SAM) and the re-training module reduces the parameter scale to the mainstream segmenters. The results validate that the re-training module contributes to an overall performance boost for semi-supervised segmentation.
@@ VIII-E Effectiveness of Re-training Module
VIII-F Visual analysis
VIII-F1 Visualization on Moderate Annotations
The single target segmentation examples on LA dataset are shown in Figure 8, compared with current SOTA methods including MT [13], UA-MT [12], DTC [43], BCP [10], MCCauSSL [17] and ACMT [9]. Our CFR and CFR predictions align more accurately with ground truth masks, further validating the effectiveness. For multiple target datasets, we demonstrate the segmentation results of MACT dataset with 14 annotated samples in Figure 10. Our CFR and CFR are compared with MT [13], UA-MT [12], CPS [15], ICT [44], SS-Net [53] and MagicNet [8] methods. It can be observed that our framework performs well in areas where semi-supervised methods struggle to identify, such as the stomach (magenta), pancreas (blue), and left kidney (green) regions.
VIII-F2 Visualization on Scarce Annotations
Figure 9 and 11 are the segmentation examples of single target and multiple target datasets with scarce annotations, taking the BraTS dataset and BTCV dataset as examples. The effectiveness of our proposed framework can be shown in some challenging examples. In these examples, the baselines (i.e., MT [13], ACMT [9] and MagicNet [8]) tend to generate false predictions or incomplete structures, whereas the introduction of the CFR framework can mitigate these problems and obtain a more plausible segmentation result.
VIII-G Further Analysis
VIII-G1 Analysis of Initial Pseudo-labels
@@
We compare our fine-tuning module of CFR framework with four foundation model fine-tuning methods, including MedSAM-v1 [19], MedSAM-v2 [19], SAMed [6] and SAMUS [20]. MedSAM adopts a sub-parts fine-tuning approach and SAMed is a LoRA-based adapter tuning method, which maintains the same parameter scale as the original SAM model. SAMUS introduces a parallel CNN branch and thus has more parameters. As shown in Figure 1, it demonstrates that our concatenation fine-tuning strategy (denoted as CFR) achieves a significant performance improvement without introducing additional parameters, and therefore could provide more reliable pseudo-labels for subsequent retraining module.
VIII-G2 Compared with 3D-based SAM methods
We compare the results of our fine-tuning CFR and the existing 3D-based SAM fine-tuning method (3DSAM-Adapter) in Table X, which reveals that our method achieves better accuracy without additional computational overhead.
VIII-G3 Compatibility under Different Number of Labeled Data
@@
With its plug-and-play nature, our re-training module could apply different SSL methods. We conduct re-training experiments on self-training, MT, and ACMT methods with different numbers of labeled samples in Figure 12. Compared to the baseline, our CFR yields excellent segmentation performance with 1, 2, 4, or 16 labeled samples. Especially in the context of extreme annotation conditions, i.e., 1 labeled sample, our framework produces impressive improvements. The comparisons with other semi-supervised methods are provided in Figure 13. The results highlight the advantages of our CFR framework in leveraging limited labeled data to achieve remarkable segmentation performance.
@@
| CFR | 91M | 90.39 | 82.54 | 1.83 | 6.71 |
VIII-G4 Spatial Structure of Medical Images
@@
During fine-tuning, regarding the distribution gap and dimensionality mismatch, we believe the global scope of all the consecutive slices could be a bridge to link the SAM (designed for 2D natural images) and 3D medical images. As we demonstrated in Section IV, our concatenation method outperforms the listed alternatives, indicating that in the fine-tuning step, SAM benefits from accurately locating the regions to segment due to its generalization ability in various segmentation tasks. During re-training, we still use 3D structure-based SSL methods, to retain spatial information. In a nutshell, SAM in fine-tuning and SSL in re-training separately treat 3D information in different ways: SAM to solve concatenated large-sized 2D slices, and SSL to further segment the 3D volumes.
IX Conclusion
@@
In this work, we present the CFR framework, which consists of the concatenate, fine-tuning, and re-training modules, to achieve higher improvements in semi-supervised segmentation tasks by leveraging the foundation model. The concatenate module copes with the resolution difference between medical and natural images, and fine-tuning module provides reliable initial pseudo-labels for re-training module. Our framework maintains the same parameter size as the mainstream segmenter, e.g., V-Net [26], and could be compatible with most popular SSL methods, e.g., Mean Teacher [13]. Extensive experiments demonstrate the CFR framework improves the performance remarkably in both moderate annotation and scarce annotation scenarios.

### v2

@@ IV-E1 Single Target Segmentation
We evaluate the impact of SFR with scarce annotations on the LA and BraTS datasets. In Table VII, when only one labeled volume is available, supervised learning performs poorly, achieving only 17.99% accuracy. SFR takes advantage of medical image stitching and SAM’s powerful feature capture capabilities to guide MT to increase from 29.68% to 74.40%. The results of the BraTS are shown in Table VIII and SFR helps MT improve by 15.06% and MagicNet by 15.15%. In addition, one phenomenon observed is that some semi-supervised methods perform worse than the lower bound. This could be attributed to the incomplete similarity of the data distribution between labeled and unlabeled samples, and the inaccurate pseudo-labels from the unlabeled samples have a negative impact on the model’s learning process. Furthermore, our SFR+ framework exploits a confidence-based selective strategy and achieves better results than SFR on both MT and ACMT methods.
We recognize that there is a large performance difference between MT with moderate and scarce annotations on the LA dataset compared to the BraTS dataset. To better understand this performance improvement with the increase of labeled samples, we perform kernel dense estimation on different settings, as shown in Fig. 10. In the LA dataset, the results show that with scarce annotation, the distribution gap between labeled and unlabeled data is quite large. This may cause MT to struggle in feature extraction because it relies on this labeled set to learn the overall data distribution. As the number of labeled samples increases, the distribution gap between labeled and unlabeled data becomes small, so the effect of MT improves rapidly. For BraTS dataset, the difference between the feature distributions of labeled and unlabeled data is not as pronounced as in LA dataset with scarce annotation, which may lead to a better effect on BraTS than on LA dataset.
IV-E2 Multiple Target Segmentation
@@
The results of scarce annotations on three multi-organ datasets are presented in Table IX, Table X, and Table XI. On BTCV dataset with only one labeled sample, the model struggles to identify esophagus, aorta, inferior vena cava, portal and splenic veins, left and right adrenal glands, with Dice scores below 5% for left kidney, stomach, and pancreas. Semi-supervised methods with unlabeled samples alleviate the poor performance of most organs. Our SFR framework achieves improvements of 10.12% and 13.57% on MT and MagicNet models, respectively, particularly with a nearly 30% improvement on aorta region. Similarly, on MACT dataset, SFR helps identify challenging classes, with significant improvements of 37.33% for spleen and 34.53% for right kidney on the Mean Teacher model. Also, SFR achieves a performance improvement of up to 10.13% (64.32% vs. 66.16%) Dice score on the AbdomenCT-1K dataset with one annotation. Our SFR+ framework further improves performance, achieving more than 10% improvement in the Dice metric on all three datasets.
Fig. 11 shows the segmentation examples of single target dataset BraTS and multiple target dataset BTCV with scarce annotation. The effectiveness of our proposed framework can be shown in some challenging examples. In these examples, the baselines (i.e., MT [21], ACMT [17] and MagicNet [16]) tend to generate false predictions or incomplete structures, whereas the introduction of SFR framework can mitigate these problems and obtain a more plausible segmentation result. Our framework demonstrates its ability to adapt flexibly to different task requirements. This versatility highlights the potential of our framework to handle a diverse range of medical imaging tasks with varying levels of annotation availability.
@@
IV-F Effectiveness of Re-training Module
We report the results of fine-tuning step omitting re-training in all datasets in Table XII, and compare the model parameters and performance after fine-tuning and re-training steps. By comparing fine-tuning results with re-training results in more detail, we observe that there is a slight performance decrease in SFR with scarce labels on LA and BTCV datasets, but overall there is a clear improvement in effect. On the other hand, considering the practical deployment requirements, re-training model has a smaller number of parameters than fine-tuning model. Our fine-tuning module SFR has the same parameter size as the foundation model (SAM) and the re-training module reduces the parameter scale to the mainstream segmenters. The results validate that the re-training module contributes to an overall performance boost for semi-supervised segmentation.
IV-G Further Analysis
IV-G1 Analysis of Initial Pseudo-labels
@@
We compare our fine-tuning module of SFR framework with four foundation model fine-tuning methods, including MedSAM-v1 [8], MedSAM-v2 [8], SAMed [10] and SAMUS [11]. MedSAM adopts a sub-parts fine-tuning approach and SAMed is a LoRA-based adapter tuning method, which maintains the same parameter scale as the original SAM model. SAMUS introduces a parallel CNN branch and thus has more parameters. As shown in Fig. 1, it demonstrates that our stitching fine-tuning strategy (denoted as SFR) achieves a significant performance improvement without introducing additional parameters, and therefore could provide more reliable pseudo-labels for subsequent retraining module.
IV-G2 Analysis of Pseudo-label Error Propagation
In our work, we use fine-tuned foundation model to generate pseudo-labels, and the errors in pseudo-labels on unlabeled data could propagate and impact the model’s performance. We have considered the potential challenge and implemented some measures to address it effectively. Firstly, within our SFR framework, the semi-supervised retraining module does not solely rely on the pseudo-labels generated by the fine-tuned SAM models. Instead, we incorporate a semi-supervised consistency loss that encourages stable predictions across various perturbations of the input. This helps the model learn consistent and robust features, reducing its sensitivity to minor inaccuracies or noise in the pseudo-labels. Secondly, in the extended version of our framework, SFR+, we introduce confidence estimation to better handle each unlabeled sample. This approach ensures that the model’s learning process is guided primarily by more reliable pseudo-labels, reducing the impact of erroneous labels on the overall performance.
IV-G3 Compatibility under Different Number of Labeled Data
@@
With its plug-and-play nature, our re-training module could apply different SSL methods. We conduct re-training experiments on self-training, MT, and ACMT methods with different numbers of labeled samples in Fig. 12. Compared to the baseline, our SFR yields excellent segmentation performance with 1, 2, 4, or 16 labeled samples. Especially in the context of extreme annotation conditions, i.e., 1 labeled sample, our framework produces impressive improvements.
@@
| SFR | 91M | 90.39 | 82.54 | 1.83 | 6.71 |
| Method | Prompt | Number | Dice | Jaccard | ASD | 95HD |
| SAM-Med3D [38] | click | 2 | 57.33 | 42.31 | 8.58 | 28.88 |
| click | 5 | 65.12 | 50.29 | 10.65 | 31.17 | |
| click | 10 | 67.17 | 52.54 | 11.05 | 33.37 | |
| SFR (Ours) - scarce annotations (1) | 76.76 | 64.65 | 8.54 | 28.07 | ||
| SFR (Ours) - moderate annotations (20%) | 90.95 | 83.47 | 1.43 | 6.11 | ||
IV-G4 Compared with 3D-based SAM methods
We compare the results of our fine-tuning SFR and the existing 3D-based SAM fine-tuning method (3DSAM-Adapter) in Table XIII, which reveals that our method achieves better accuracy without additional computational overhead. Our method outperforms the 3DSAM-Adapter by 8%. A possible explanation for this performance difference lies in the size of the training parameters. For small-scale medical imaging datasets, particularly in scenarios where training data is limited, a model with a larger number of parameters, which has high complexity, may be prone to overfitting. In contrast, our method effectively utilizes 2D slice stitching input with the low-rank-based fine-tuning strategy. It allows the model to learn spatial relationships while managing computational complexity, which results in better generalization and segmentation performance.
In addition, we conduct a comparison with the SAM-Med3D method [38], which modifies the SAM paradigm to 3D architecture with training from scratch on 3D medical image datasets. We compare the results of our framework with the prompt-based SAM-Med3D method on the same test data in Table XIV. Notably, on the LA dataset, our SFR framework, using just a single annotation, outperforms SAM-Med3D with 2, 5, or 10-point prompts per sample.
IV-G5 Compared with SAM 2-based methods
We evaluated both SAM 2 [31] and Medical SAM 2 [34] across multiple datasets, including LA, BraTS, BTCV, and MACT, under different prompt conditions (2, 5, and 10 clicks), as shown in Table XV. While these models perform well in certain datasets, such as BTCV and MACT, our SFR framework demonstrates significantly superior performance when moderate annotations are provided. Moreover, for the LA and BraTS datasets, the Dice scores of our method, even with scarce annotations (1 annotation), are noticeably higher than the results obtained by both SAM 2 and Medical SAM 2 under any prompt frequency condition. This further underscores the robustness of our approach, particularly when dealing with limited annotations.
| Method | Prompt | Frequency | LA | BraTS | BTCV | MACT |
| SAM 2 [31] | click | 2 | 31.13 | 25.68 | 55.64 | 77.09 |
| click | 5 | 30.73 | 25.08 | 54.07 | 73.12 | |
| click | 10 | 29.41 | 23.88 | 53.48 | 72.40 | |
| Medical SAM 2 [34] | click | 2 | 31.64 | 28.91 | 59.47 | 78.77 |
| click | 5 | 33.56 | 28.06 | 56.78 | 76.67 | |
| click | 10 | 30.38 | 25.85 | 56.27 | 75.42 | |
| SFR (Ours) - scarce annotations | 76.76 | 78.58 | 53.59 | 49.08 | ||
| SFR (Ours) - moderate annotations | 90.95 | 85.81 | 77.06 | 82.87 | ||
IV-G6 Compared with Concatenating along Channel Dimension
By concatenating three slices along the channel dimension to create a pseudo-RGB image, SAM could potentially segment this combined representation. On the other hand, as the SAM model is pre-trained on natural color images, it might interpret the concatenated channels more as color information than as meaningful anatomical context. This presents an intriguing yet challenging question, that is worth exploring further in future research to better evaluate its applicability and uncover potential benefits.
IV-G7 Spatial Structure of Medical Images
@@
During fine-tuning, regarding the distribution gap and dimensionality mismatch, we believe the global scope of all the consecutive slices could be a bridge to link the SAM (designed for 2D natural images) and 3D medical images. As we demonstrated in Section III-B1, our stitching method outperforms the listed alternatives, indicating that in the fine-tuning step, SAM benefits from accurately locating the regions to segment due to its generalization ability in various segmentation tasks. During re-training, we still use 3D structure-based SSL methods, to retain spatial information. In a nutshell, SAM in fine-tuning and SSL in re-training separately treat 3D information in different ways: SAM to solve stitched large-sized 2D slices, and SSL to further segment the 3D volumes.
V Conclusion
@@
In this work, we present the SFR framework, which consists of the stitching, fine-tuning, and re-training modules, to achieve higher improvements in semi-supervised segmentation tasks by leveraging the foundation model. The stitching module copes with the resolution difference between medical and natural images, and fine-tuning module provides reliable initial pseudo-labels for re-training module. Our framework maintains the same parameter size as the mainstream segmenter, e.g., V-Net [30], and could be compatible with most popular SSL methods, e.g., Mean Teacher [21]. In addition, we develop the SFR+, which further enhances the framework by introducing confidence estimation and selective training strategy. Extensive experiments demonstrate that the SFR and SFR+ frameworks improve performance remarkably in both moderate and scarce annotation scenarios.